OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 3x + y = 26 -6x + y = -37

OpenStudy (anonymous):

3x + y = 26 y= 26 - 3x -6x + (26 - 3x) = -37 -9x + 26 = -37 -9x = -63 x = 7 3(7) + y = 26 y = 26- 21= 5

OpenStudy (anonymous):

Wow! Thank you. I am very grateful. I am having some trouble with this course.

OpenStudy (anonymous):

you're welcome! If you don't understand anything else, just ask.

OpenStudy (anonymous):

Could you possible show me step by step on this one? The tutors in my school don't even get them correct. Lol. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 10x + 5y = 52.5 30x – 7y = 80.5

OpenStudy (anonymous):

possibly*

OpenStudy (anonymous):

Elimination is when you subtract or add one equation from another equation to get rid of one of their variables and be able to solve for the other. Does that make sense?

OpenStudy (anonymous):

Yes. It does. I just bog down when I see the questions .

OpenStudy (anonymous):

If we're going to eliminate x to solve for y, first we have to do this: Multiple both sides of 10x + 5y = 52.5 by 3 3(10x + 5y) = 3(52.5) 30x + 15y=157.5 Now, we subtract this equation from the other 30x + 15y=157.5 - 30x – 7y = 80.5 --------------------- 0 + 22y=238 y=238/22

OpenStudy (anonymous):

Do all of those steps make sense to you? And can you solve for x now?

OpenStudy (anonymous):

I think I got it. Thank you! I am sorry, I have to go. You were a big help!! :)

OpenStudy (anonymous):

I'm glad!